17. Phosgene, substance used in poisonous gas warfare a during World War I, is so named because it was first on a mixture of carbon prepared by the action of sunlight from the Greek monoxide and chlorine gases. Its name comes Words phos (light) and genes (born of). Phosgene has the
following elemental composition: 12.14% C, 16.17% O, 71.69% and Ci by mass. is 98.9 g/mol. 10pt/ea this compound. Its molar mass a. Determine the molecular formula of b. Draw three Lewis structures for the molecule that (The C1 and O satisfy the octet rule for each atom. atoms bond to C.) Using formal charges, determine which Lewis structure is the dominant one. C. the molecule and write its d. Sketch the 3D shape for molecular geometry name.
Answer #1

Let there is 100g of the phosgene.

Number of moles of Carbon = 12.14g/12g/mol = 1.012

Number of moles of Oxygen = 16.17g/16g/mol = 1.011

Number of moles of chlorine = 71.69g/35.45g/mol = 2.022

Simple molar ratio of C:O:Cl = 1.012:1.011:2.022

= 1:1:2

(A) molecular formula of the compound = COCl2

(B)

11 :Ce CO: ce: Roamal change thi uts fas no and each element fos comblede octet Hence Steble lewis svcu Ce Triangular planar